Susan Marciano invested part of $34,000 bonus in a fund that paid a 12% profit and invested the rest in stock that suffered a 5% loss. Find the amount of each investment in her overall net profit was $170.
The amount invested at 12 % is=
The amount invested in stock is =
2. S= 4LW+ 2WH; S= 84, L= 9, W=2
H=______
3. Country A produces about 9 times the amount of diamonds in carats produced in country B. If the total produced in both country is 20,000,000 carats find the amount produced in each country.
Country A=
Country B=

Answer by Boreal(15235)

x and 34,000-x
.12x-.05(34000-x)=170
.12x-1700+.05x=170
.17x=1870
17x=187000
x=$11000@12%=$1320
32000-x@=$23000@-5%=-$1150
The gain is $170.
$11,000 in fund
$23,000 in stocks.
------------------------
S= 4LW+ 2WH; S= 84, L= 9, W=2
84=4*9*2+4W=72+4H
12=4H
H=3
-----------------------
A=9x
B=x
Their sum is 10x=20,000,000
x=2 million
9x=18 million, which is 9 times as much and the two add to 20 million.